Required information An air-standard cycle is executed within a closed piston, cylinder system, and it consists of the followAn air-standard cycle is executed within a closed piston–cylinder system, and it consists of the following three processes: 1–2 V = Constant heat addition from 100 kPa and 30°C to 850 kPa 2–3 Isothermal expansion until V3 = 8.5V2 3–1 P = Constant heat rejection to the initial state Assume air has constant properties with cv = 0.718 kJ/kg·K, cp = 1.005 kJ/kg·K, R = 0.287 kJ/kg·K, and k = 1.4.
Answer #1

In this problem, first draw P-V diagram for the given cycle then use area under the curve for Calculation of thermal efficiency which is Calculated as below:-

Given data :- *T=constant 100 ka P, = 100 kPa Ti= 30°c= 303 K 1850k P₂ = 850 kla Vg = V2(8.5) for air, Co=0.718 kolkg_k cp =3- 4- therma u Wan & Area (2-3-4-5-1) = I go thermal proces work - Peva lim ( 13 ) = Pq Vz ln (8.582) = Pe V2 ln (8.5) - 2.14